OpenStudy (anonymous):

The attendance on the first day of a carnival was 425 people. The attendance on the third day was 575 people. Assuming attendance will increase linearly each day, how many people will attend the carnival on the sixth day? 650 725 800 875

OpenStudy (l):

575-425 = 150 people in 3 days 150/3 = 50 people in a day 50*6 = 300 people in 6 days 425 + 300 = 725 people on the sixth day *

OpenStudy (anonymous):

thanks so much, your AMAZING (: